OpenStudy (anonymous):

3 squareroot of 2 -5 squareroot of 8 please help!

OpenStudy (anonymous):

okay, first you need to simplify \[-5\sqrt{8}\] do you know how to do that?

OpenStudy (anonymous):

@kmallum u there?

OpenStudy (anonymous):

dont remember! @mkerr017

OpenStudy (anonymous):

thats fine, you just seperate it into \[-5\sqrt{2\times4}\] then, since four is a perfect square, you can take it out, by converting it to a 2. Multiply the -5 by the 2 that you took out and you get \[-10\sqrt{2}\]

OpenStudy (anonymous):

now we are down to \[3\sqrt{2} -10\sqrt{2}\] since they have the same radical, you can combine the two outside numbers and you end up with \[-7\]

OpenStudy (anonymous):

sorry, \[-7\sqrt{2}\]

OpenStudy (anonymous):

ohh i see thank you! @mkerr017

OpenStudy (anonymous):

no problem:) let me know if you have any more questions

OpenStudy (anonymous):

how about, 2 squareroot45 - squareroot 80 + 3 squareroot 20

OpenStudy (anonymous):

do you want explanation, or just answer?

OpenStudy (anonymous):

explaination.. im just a bit off on the coefficient .. the base is squareroot 5.. and thats correct.. the answer is supposed to be 8 squareroot 5

OpenStudy (anonymous):

K, so the first thing to look at is that the radical for all of these is different and we need to get them to be the same. By looking i can see that i can get them down to 5. I'll take 9 out of the first one, 16 out of the second, and 4 out of the third leaving me with \[6\sqrt{5}\]\[-4\sqrt{5}\] and \[6\sqrt{5}\] then I simply combine them to get \[8\sqrt{5}\] like you said:)

OpenStudy (anonymous):

( 5 squareroot of 2) ^2 - (6 squareroot of 3) ^2 ? :)

OpenStudy (anonymous):

sorry, i don't know how to do this one:(